# Code Prompting Elicits Conditional Reasoning Abilities in Text+Code LLMs

### 18 Jan 2024 v1

@@ 5 Experiments
We first compare the performance of the two prompting methods — text prompts and code prompts. We then conduct extensive ablation experiments to determine whether it is the code-translated instances what improve performance or if the improvement is caused by implicit text simplification (Section 5.2) or by models merely being exposed to code within prompts and not necessarily the instances translated to code (Section 5.3). Finally, we show that code prompting is more sample efficient (Section 5.4) when compared to text prompting and that models prompted with code exhibit superior state tracking capabilities (Section 5.5).
@@ 5.1 Code Prompting Improves over Text Prompting
Overall, as shown in Table 1, code prompts consistently outperform text prompts. In particular, we can observe a significant performance boost of $7.71$ points on BGQA-3, $7.4$ on BGQA-2, $6.6$ on BGQA-1, and $2.6$ points on CondQA. Furthermore, we can also observe how the gap between both prompting methods increases in BGQA as the difficulty of the task increases. This clearly indicates the potential benefits of using code to elicit reasoning abilities, especially on the multi-hop subsets.
@@
We note that these consistent and substantial gains using code prompts are obtained despite the straightforward transformation of the text prompt, which does not incorporate new information as shown in Figure 2. This suggests that the code format exhibits characteristics that trigger the conditional reasoning abilities of text+code LLMs.
In the following experiments, we perform detailed ablation studies confirming that the performance improvements indeed originate from the code format and render additional insights into the potential reasons behind this phenomenon.
Prompt CondQA CondQA-YN BGQA-1 BGQA-2 BGQA-3
Code Prompt $\mathbf{60.04}$ $\mathbf{78.94}$ $\mathbf{70.2}$ $\mathbf{63.2}$ $\mathbf{55.0}$
Anonymous code $58.42$ $76.02$ $63.6$ $58.4$ $51.0$
Random code $56.64$ $76.27$ $62.8$ $54.0$ $45.2$
- Comments N.A. $64.92$ $53.5$ $47.0$ $49.8$
Table 2: Ablation study on the relevance of code semantics within Code Prompts. CondQA-YN represents the partition of ConditionalQA with yes-no answers.
5.2 Code Format Elicits Reasoning Abilities
We are now interested in whether the observed improvements in code prompting originate from the input instances being formatted as code or if they are caused by the implicit simplification of text that occurs when the instances are translated into code. To investigate this, we devise experiments with prompts that represent the intermediate states between natural language and code.
@@ 5.2 Code Format Elicits Reasoning Abilities
Inspired by Min et al. (2023), in our first study, we transform each NL sentence [^3] We only transform the facts in BGQA since transforming the rules into atomic statements as well yields worse results. into a sequence of atomic statements, which we then append to the original sentence. In this way, the atomic statements can be seen as defining variables for each key entity in the text. Therefore, this new type of prompt would resemble code but without control flow and in natural language form. For instance, for the sentence Applying for the legal right to deal with someone’s property, money and possessions (their estate) when they die is called applying for probate, some atomic statements would be: i) Applying for the legal right is a process and ii) The process is called ‘applying for probate’. A complete example is provided in Appendix D.
The prompt retains access to the original instance text (i.e., no loss of information) but is also augmented by simplified sentences in the form of atomic statements. This setup allows us to investigate whether the simplicity of the input prompt is the cause of improved reasoning abilities, regardless of the text and code formats.
@@
In our second experiment, we investigate whether the semantics of the code statements and not the form of the code are the reason behind the performance boost. For this purpose, we back-transform the code prompts into NL such that the reasoning statements (i.e., the if conditions) are clearly and concisely stated in natural language. Specifically, we map every variable into the format Key entity: variable without snake case. For instance, the variable husband_pass_away from Figure 2 would be back-transformed as Key entity: husband pass away. To transform the if statements, we create a translation prompt by providing four demonstrations. These demonstrations simply translate the conditional statements within the  code-formatted instance back into natural language in the same manner as the variables. This makes the back-translated text as close as possible to the code text. We provide examples of this in Table 6 from Appendix E.
From the results in Table 1, we see that (1) prompting with atomic statements does not yield performance improvements w.r.t. text prompts, and (2) mapping back from code to NL provokes performance drop w.r.t. code prompts, yielding similar results to the original text prompt. In the first case, we observe that the performance drop w.r.t. the original text prompt can be due to the sampling variance since the results are within the standard deviation of text prompts. However, we obtain a more acute drop for BGQA. This can be attributed to the format of this dataset. The atomic statements method we use (Min et al., 2023) was devised for general texts such as Wikipedia pages. However, BGQA is a logic-based dataset, where the input facts are statements that are already minimal units of information and do not follow the style of general documents. Thus, creating atomic statements from these sentences can break the sentence structure needed to track the attributes of the subjects and objects of the sentences.
These results indicate that the enhanced complex reasoning abilities do not originate from input text simplification nor the conditional statements being extracted from the instance sentences, but rather the code format of the input instances.
5.3 Code Semantic is Important
@@
Previously, we have shown that the code format is necessary to elicit the reasoning abilities of text+code LLMs. Now, we aim to investigate which aspects of code are needed for this. In particular, we evaluate the impact of retaining the natural language text of the original instance within the code comments and the importance of the semantics of the code. To analyze the former, we have removed the code comments that include the original natural language text from the input and evaluated the performance of the new prompts. To analyze the latter, we perturbed the code to anonymize the variables and functions, as well as added random code whose semantics are completely irrelevant to the original natural language text. In the latter two cases, the code comments remain unmodified (examples illustrating them are provided in Table 7 from Appendix E). Since CondQA includes span answers and removing the NL text would make it impossible for the model to generate the span, we only report performance on the yes-no answers partition (i.e., CondQA-YN).
@@
Code perturbations (anonymous code and random code) also confirm the importance of code semantics to elicit reasoning abilities. When we use anonymized code, we observe a performance reduction of almost $2$ points on CondQA and a decrease between $6.6$ and $4$ in BGQA. The decrease is even larger when the code is randomized, with drops of more than $3$ points on CondQA and between $7.4$ and $9.8$ on BGQA. This more pronounced drop is expected since the semantics and logic of the code mismatch the NL text, whereas anonymous code maintains the same logic on both NL and code.
Furthermore, we also observe that the performance of random code prompts is similar to that of text prompts (Table 1) in all datasets. This can be interpreted as the model being able to identify the irrelevance of the code to the text. Hence, the model disregards the code to solely focus on the code comments (i.e., the natural language text). This could be possible thanks to the provided demonstrations, which show answers that only refer to the natural language text.
These results confirm that code alone does not trigger reasoning abilities, and instead, the combination of code that represents the original natural language instance and the NL text exploits the potential of LLMs.
@@ 5.4 Code Prompts are More Sample-Efficient at Eliciting Reasoning Abilities
Given our observations that code prompts trigger conditional reasoning abilities better than text prompts, it is natural to ask the follow-up question: are code prompts also more sample-efficient at triggering conditional reasoning abilities when compared to text prompts? To answer this, we evaluate how the overall performance of the model changes with respect to the number of demonstrations.
In the following experiments, we compare the performance gap of the two prompting methods under different numbers of demonstrations. Figure 3 shows that when we only provide one demonstration per class (i.e., answer type in our datasets), the performance gap is the largest across all datasets. As expected, this gap decreases when we provide more demonstrations. These results indicate that code prompts trigger conditional reasoning more efficiently than text prompts, and this is one of the reasons for its superior performance. This phenomenon has further applicability in long-document and very low resource settings, where only one demonstration is available or fits within the input window, and therefore, it would be advisable to convert that demonstration into code.
Figure 3: Performance comparison between text prompts (blue) and code prompts (green) using 1, 2, and 3 demonstrations per class.
@@ 5.5 Code Prompts Improve Variable Tracking in LLMs
Correct Instances Incorrect Instances
Dataset Text Prompts Code Prompts Text Prompts Code Prompts
@@
Table 3: Comparison of the percentage of memory errors made by GPT 3.5. For each dataset, we separately compute memory errors for the instances where the model gives the correct and incorrect answers. Lower is better.
We hypothesize that one of the reasons for the superior performance of code prompting is an improved ability to track the states of variables or key concepts. This hypothesis is based on the intuition that, for natural language in general, local context is the most important part of generating the next token Khandelwal et al. (2018); Sun et al. (2021). However, generating code is often more challenging because code frequently refers to previously defined functions, which can be dozens or even hundreds of lines apart. This resembles multi-hop reasoning, where the model may need to reference a key entity dozens of lines before. Therefore, an improved ability to look for distant co-references can be beneficial to multi-hop reasoning, which is also needed to solve our datasets.
[…]
@@
We acknowledge that the output text may not be completely faithful to the internal beliefs of the model (Lyu et al., 2023). Therefore, we first test whether this experiment can be a proxy metric of the internal belief of the model. To do this, we compare the memory error percentage of the prompting methods in instances where the model solves (i.e., correct instances) and does not solve (i.e., incorrect instances) the question. If incorrect instances yield a higher memory error, this would indicate that the model struggles more to remember the state of those variables, and this would make it fail to conduct the reasoning process. Therefore, our probes would be a proxy metric of the internal belief of the model.
@@
Table 3 shows the results of this comparison. We observe that all prompting methods in all datasets consistently make more memory mistakes on incorrect instances than on correct instances, with the exception of text prompts on CondQA. However, the memory error in this case is significantly high, which may suggest that the model is not able to track entities correctly in any case. Therefore, we can use this experiment as a proxy measure of the memory of the model.
@@
From the same table, we can observe that Text Prompts make significantly more memory errors than code prompts on all datasets. Specifically, the gap is consistently more than 30% with peaks on CondQA (66.69%) and BGQA-3 (39.8%). Therefore, this experiment empirically confirms our hypothesis that code prompts can track the state of the variables/key entities better than text prompts.
5.6 Qualitative Analysis
All our previous experiments have focused on the accuracy of the answers. Obtaining a correct answer is correlated to a correct reasoning process. However, it is possible to obtain a correct answer despite an incorrect reasoning chain. Due to the difficulty of automatically evaluating the chain of thoughts, we perform a limited human evaluation. We sample ten instances from BGQA-3 (i.e., the dataset with the most complex reasoning chains) where both text and code prompts return the correct answer and manually inspect the quality of the reasoning chains. Based on this limited sample, we observe that text prompts conduct a 100% correct reasoning chain in only two cases, while code prompts do so in three cases. We identify two main types of errors: (1) wrong conditional reasoning and (2) commonsense errors. We provide examples of those in Table 4. This observation raises the question of whether the generated chain of thought is not faithful to the internal reasoning of the model, as suggested by (Lyu et al., 2023) or whether the model generated the right answer from a greedy attempt to reach the closest plausible conclusion (code prompts shows a reasoning error in the last step of the CoT in three out of seven cases). We leave the analysis of faithfulness of chains of thoughts as future work.
Error Type Example Explanation
Commonsense
The model believes a harmonica is something to sit on.
Conditional Reasoning
become_enemy(lion, squid)==True
not remove_piece(lion, dog)==True
steal_points(lion,mosquito,5)=rule5(lion)
steal_points(lion, mosquito, 5)==True
@@ 6 Conclusions and Future Work
This work demonstrates that conditional reasoning abilities can be triggered by using code prompts in large language models (LLMs) of text and code. These code prompts contain the original natural language (NL) formulation as a code comment and code that formulates the logic of the text (e.g., documents and questions). To create these code prompts, we use in-context learning to teach an LLM how to automatically conduct such a transformation. Through multiple experiments, we show that code prompts trigger conditional reasoning abilities. They consistently and substantially outperform text prompts across two conditional reasoning datasets by between $2.6$ and $7.7$ points. Our experiments show that even simple code can be beneficial as long as it closely follows the semantics of the NL text and is accompanied by the original NL text. In addition, code prompts require fewer demonstrations, making them more efficient than vanilla text prompts. Our experiments also suggest that code prompts achieve the performance boost due to their superior ability to track the state of variables or key entities.
In our future work, we plan to extend our experiments to other text+code LLMs. We also plan to conduct experiments investigating how pretraining on text, code, and text+code affects the triggering of reasoning abilities. Lastly, we would like to extend our experiments to a wider range of reasoning abilities to verify the generalization of our findings.
@@
Transforming a natural language problem into code requires an intermediate step that raises the cost of the whole pipeline. However, this mapping is not a complicated task. Therefore, we believe it would be possible to train a small generative model to do it instead of using a large language model. In this way, we could minimize the cost of using code prompts without affecting its performance.
Due to the costs of very large language models and the large size of BoardgameQA, we only ran the experiments two times with different random seeds. However, the variance is small enough to ensure the representativeness of the results.

### 25 Feb 2024 v2

@@ 5 Experiments
[…]
@@ 5.1 Code Prompting Improves over Text Prompting
Table 1 shows the model performance on the development and test sets. Code prompts outperform text prompts in the majority of cases on the test set (11 out of 15). This trend holds true across models, with each achieving peak performance through code prompts for most datasets (i.e., GPT-3.5 in 4/5, Mixtral in 3/5, Mistral in 4/5). Notably, code prompts consistently surpass text prompts on BGQA-2 and BGQA-3, the most reasoning-intensive datasets (see Appendix A), for all models. This is particularly evident for GPT-3.5, where gains exceed 18 points. Conversely, the advantage is narrower on CondQA, where the linguistic dimension plays the biggest role (see Appendix A). This suggests that code prompts elicit conditional reasoning abilities and are most suited for reasoning-intensive tasks. Furthermore, in the cases where text prompts are superior, their average gains are only 3.29. In contrast, code prompts lead to significantly larger mean and median gains of 10.48 and 7.75, respectively, in the cases where they are superior. Additionally, an experiment with Phi-2, a small language model, reveals a substantial 15-point performance improvement using code prompts (see Appendix E).
To shed light on the performance gains driven by code prompts, we delve into the confusion matrices (attached in Appendix J) and discover that text prompts in Mistral predict “not enough information” much less than code prompts for BGQA. This is particularly noticeable in BGQA-1, where text prompts do not predict a single “not enough information,” while code prompts do. On the other hand, text prompts in GPT 3.5 and Mixtral overpredict “not enough information” on BGQA and ShARC, leading to a low number of true positives for the conclusive answers. We hypothesize that this model hesitation could stem from the alignment tax (Ouyang et al., 2022) of reinforcement learning from human feedback models. This potential barrier may be alleviated by code prompts because they indicate to the model the variable that answers the question and instruct the model to track the entailment status of variables within the given code.
@@
These consistent and substantial gains from code prompts are obtained despite a straightforward transformation of text prompts, which does not incorporate new information, as shown in Figure 2. This finding strongly suggests that code possesses specific characteristics that effectively elicit conditional reasoning abilities within text+code LLMs.
5.2 Code Syntax Elicits Reasoning Abilities
We now want to delve into the source of the performance gains observed when using code prompting. We investigate whether these improvements stem from the simplification of text into premises facilitated by code, effectively reducing the task to a form of semantic inference within the linguistic dimension, or if there are inherent properties of code syntax that contribute to enhanced performance. To investigate this, we devise experiments with prompts that represent the intermediate states between natural language and code.
@@ 5.2 Code Syntax Elicits Reasoning Abilities
I. Atomic Statements.   Inspired by Min et al. (2023), we transform each NL sentence [^5] We only transform the facts in BGQA since transforming the rules into atomic statements as well yields worse results. into a sequence of atomic statements, which we then append to the original sentence. In this way, the atomic statements can be seen as defining variables for each key entity in the text. Hence, this new prompt would resemble code but without control flow and in natural language form. The prompt retains access to the original instance text (i.e., no loss of information) but is also augmented by simplified sentences in the form of atomic statements. This setup allows us to investigate whether the simplicity of the input triggers improves reasoning abilities, regardless of the text and code syntax.
@@
II. Back-Translated Code.   In our second experiment, we investigate whether the semantics of the code statements and not the code syntax are the reason behind the performance boost. For this purpose, we back-transform the code prompts into NL such that the reasoning statements (i.e., the if conditions) are clearly and concisely stated in natural language. Specifically, we map every variable into the format Key entity: variable without snake case. For instance, the variable husband_pass_away from Figure 2 would be back-transformed as Key entity: husband pass away. To transform the if statements, we create a translation prompt by providing four demonstrations. These demonstrations simply translate the conditional statements within the  code-formatted instance back into natural language. We also translate the variables in the same manner. This makes the back-translated text as close as possible to the code text. We provide examples of this in Table 11 from Appendix H.
Results.   The results [^6] We do not conduct ablation tests on ShARC because, as explained in Section 5, these ablations aim to understand why code prompts outperform text prompts using the highest performing model. in Table 2 show that (1) prompting with atomic statements does not reach the performance of code prompts, and (2) mapping back from code to NL results in a performance drop compared to code prompts. These findings suggest that code prompts enhance LLM performance beyond mere text simplification. This conclusion is supported by the observation that these alternative text simplification approaches, despite offering similar semantics to code prompts, fail to replicate the performance gains observed with code prompts. Therefore, these results imply that specific syntactic features embedded within code directly contribute to performance improvement.
Lastly, our evaluation on BGQA-3 reveals a significantly larger performance decline when using atomic statements compared to back-translated code. This disparity likely stems from the dataset’s inherent structure. The method we employ for generating atomic statements (Min et al., 2023) was specifically designed for general text formats like Wikipedia pages. However, BGQA is a logic-based dataset where input "facts" are already presented as minimally informative statements, deviating from the typical structure of general documents. As a result, generating atomic statements from these sentences can unintentionally disrupt the sentence structure, making it difficult to track the attributes of subjects and objects within the text. This observation is further supported by our results on CondQA, a dataset with longer documents, where atomic statements achieve higher performance than back-translated code.
Dataset $\Delta$ Atomic St. $\Delta$ Code normal-$\rightarrow$ NL
CondQA $-2.66$ $-4.72$
BGQA-1 $-4.37$ $-1.43$
BGQA-2 $-8.72$ $-5.39$
BGQA-3 $-19.26$ $-3.68$
Table 2: Performance gap of atomic statements and back-translated code when compared to code prompts using GPT 3.5. Results from the dev set of each dataset.
5.3 Code Semantics are Important
@@
Previously, we have shown that code syntax is necessary to elicit the reasoning abilities of text+code LLMs. Now, we aim to investigate which aspects of code are pivotal. In particular, we evaluate the impact of retaining the natural language text of the original instance within the code comments and the importance of the code semantics. To analyze the former, we have (1) removed the code comments that include the original natural language text from the input and evaluated the performance of the new prompts. To analyze the latter, we (2) perturbed the code to anonymize the variables and functions, as well as (3) added random code whose semantics are completely irrelevant to the original natural language text. In the latter two cases, the code comments remain unmodified (examples illustrating them are provided in Table 12 from Appendix H). Since CondQA includes span answers and removing the NL text would make it impossible for the model to generate the span, we only report performance on the yes-no answers partition (CondQA-YN).
@@
Effect of Code Perturbations.
Code perturbations (anonymous code and random code) confirm the importance of code semantics in eliciting reasoning abilities. When we use anonymized code, we observe a performance reduction of almost $2$ points on CondQA and a decrease between $6.6$ and $4$ in BGQA. The decrease is even larger when the code is randomized, with drops of more than $3$ points on CondQA and between $7.4$ and $9.8$ on BGQA. This more pronounced drop is expected since the semantics and logic of the code mismatch the NL text, whereas anonymous code maintains the same logic on both NL and code. Furthermore, we also observe that the performance drop of random code prompts is similar to that of text prompts (Table 1) on CondQA and BGQA-1. This can be interpreted as the model being able to identify the irrelevance of the code to the text. Hence, the model disregards the code to solely focus on the code comments (i.e., the natural language text). This could be possible thanks to the provided demonstrations, which show answers that only refer to the natural language text.
These results confirm that code alone does not trigger reasoning abilities, and instead, the combination of code that represents the original natural language instance and the NL text is able to unlock the potential of LLMs.
Prompt CQA CQA-YN BG1 BG2 BG3
Anonym. $-1.62$ $-2.90$ $-6.60$ $-4.80$ $-4.00$
Random $-3.40$ $-2.67$ $-7.40$ $-9.20$ $-9.80$
- Comments N.A. $-14.02$ $-16.70$ $-16.20$ $-5.20$
Table 3: Performance gap to code prompts for each code perturbation. cQA stands for CondQA, CQA-YN for the partition of CondQA with yes-no answers, BG for BGQA. Results reported on the dev set of each dataset.
@@ 5.4 Code Prompts are More Sample-Efficient at Eliciting Reasoning Abilities
Given our observations that code prompts trigger conditional reasoning abilities better than text prompts, it is natural to ask the follow-up question: are code prompts also more sample-efficient than text prompts? To answer this, we evaluate how the overall performance of GPT 3.5 changes with respect to the number of demonstrations for the two prompting methods.
Figure 3 shows that when we only provide one demonstration per class (i.e., answer type in our datasets), the performance gap is the largest across all datasets. As expected, this gap decreases when we provide more demonstrations. Moreover, we also observe that code prompts with only one demonstration per class even outperform text prompts with three demonstrations per class, which further shows the sample efficiency of code prompts. These results indicate that code prompts trigger conditional reasoning more efficiently than text prompts on GPT 3.5, and this is one of the reasons for its superior performance.
Figure 3: Performance comparison of GPT 3.5 between text prompts (green) and code prompts (blue) using 1, 2, and 3 demonstrations per class. Results from the dev set of each dataset.
@@ 5.5 Code Prompts Improve Variable Tracking in LLMs
We hypothesize that one of the reasons for the superior performance of code prompting is an improved ability to identify and track the states of key variables or concepts. This hypothesis is based on the intuition that, for natural language in general, local context is the most important part to generate the next token Khandelwal et al. (2018); Sun et al. (2021). However, generating code is often more challenging because code frequently refers to previously defined functions and variables, which can be dozens or even hundreds of lines apart. This resembles multi-hop reasoning, where the model may need to reference a key entity dozens of lines before. Therefore, an improved ability to look for distant co-references caused by training on code can be beneficial for multi-hop reasoning, which is also needed to solve our datasets.
[…]
Correct Ans. Incorrect Ans.
Dataset Text Code Text Code
@@
Table 4: Comparison of the percentage of memory errors made by GPT 3.5. For each dataset, we separately compute memory errors for the instances where the model gives the correct and incorrect answers. Lower is better. Results from the dev set of each dataset.
@@
Does Generated Text reflect Model Beliefs?   As the generated text may not be faithful to the internal beliefs of the model (Lyu et al., 2023), we first test the validity of this experiment as a proxy metric of the internal belief of the model. To do this, we compare the memory error percentage of the prompting methods in instances where the model solves (i.e., correct instances) and does not solve (i.e., incorrect instances) the question. If incorrect instances yield a higher memory error, this would indicate that the model struggles more to remember the variable states on those instances, which in turn would make it more likely to fail when conducting the reasoning process. Therefore, our probes would be a proxy metric of the internal belief of the model.
@@
Table 4 shows the results of this comparison. We observe that all prompting methods in all datasets consistently make more memory mistakes on incorrect instances than on correct instances, with the exception of text prompts on CondQA. However, the memory error in this case is significantly high, which may suggest that the model is not able to track entities correctly in either case. Therefore, we can use this experiment as a proxy measure of the memory of the model.
@@
Code Prompting improves State Tracking.   From Table 4, we further observe that Text Prompts make significantly more memory errors than code prompts on all datasets. Specifically, the gap is consistently more than 30% with peaks on CondQA (66.69%) and BGQA-3 (39.8%). Therefore, this experiment empirically confirms our hypothesis that code prompts improve state tracking of the key entities and variables when compared to text prompts.
@@ 6 Conclusions and Future Work
This work demonstrates that conditional reasoning abilities can be triggered by using code prompts in large language models (LLMs) of text and code. These code prompts contain the original natural language (NL) formulation as a code comment and code that formulates the logic of the text. To create these code prompts, we use in-context learning to teach an LLM how to conduct such a transformation automatically. Through multiple experiments on three LLMs and three datasets, we show that code prompts trigger conditional reasoning abilities, with large performance gains w.r.t. text prompts (up to $22.52$ percentage points on GPT 3.5, $7.75$ on Mixtral, and $16.78$ Mistral). Our experiments show that even simple code can be beneficial as long as it closely follows the semantics of the NL text and is accompanied by the original NL text. We also show that code prompts are more sample-efficient than text prompts and that their performance boost stems from their superior ability to identify and track the state of key variables or entities, a central aspect of the logical dimension of semantic inference.
In our future work, we plan to extend to a wider range of reasoning abilities, such as multi-hop reasoning, to verify the generalization of our findings. We also plan to conduct experiments investigating how pretraining on text, code, and text+code affects the triggering of reasoning abilities.
@@
Transforming a natural language problem into code requires an intermediate step that raises the cost of the whole pipeline. However, this mapping is not a complicated task, as even the smallest models we considered were able to perform it successfully. Therefore, we believe it would be possible to train a small generative model to do it instead of using a large language model. In this way, we could minimize the cost of using code prompts without affecting its performance.
We only ran the experiments on the dev set with two different random seeds due to the costs of running large language models and because we prioritized experimenting on multiple models and datasets. Nevertheless, the results of all models exhibit similar patterns, which confirms the representativeness of our results. Also, we conduct the ablations only on GPT 3.5, the best-performing and largest model. However, confirming that the findings from these ablations also hold on the smaller models would be interesting.
Lastly, we conduct our experiments on data in English. Analyzing whether our findings hold true in other languages would be interesting. However, the lack of conditional reasoning datasets in other languages would make this study difficult.
Ethics and Broader Impact Statement
It is our sincere goal to contribute to the social good in multiple ways. Firstly, we note that large language models are massively being adopted by companies to improve their products. However, it is also known their limitations, such as hallucinations and lack of faithfulness, among others. Our work aims to improve the reasoning abilities of LLMs. Also, the use of code prompts may simplify the explainability of the model responses since we can inspect the entailment status of the variables. We hope these results contribute to enhancing the trustworthiness and safety of LLMs. Nevertheless, every development may pose some risks. In our case, the improvement of the reasoning abilities in LLMs may utilized by malicious actors to propagate more persuasive disinformation.